# A Novel Technosignature Search in the Breakthrough Listen Green Bank Telescope Archive

### v1

@@ Abstract
The Breakthrough Listen program is, to date, the most extensive search for technological life beyond Earth. As part of this goal, over the past nine years it has surveyed thousands of nearby stars, close to 100 nearby galaxies, and a variety of exotic and solar system objects with telescopes around the world, including the Robert C. Byrd Green Bank Telescope (GBT) in West Virginia. The goal is to find evidence of technosignatures of other civilizations, such as narrowband Doppler drifting radio signals. Despite the GBT’s location in a radio-quiet zone, the primary challenge of this search continues to be the high quantities of human-generated radio-frequency interference (RFI), and the ability to pick out genuinely promising candidates from it. Here we present a novel search method aimed at finding these ‘needle-in-a-haystack’ type signals, applied to 9,684 observation cadences of 3,077 stars (each observed with one or more of the L, S, C, and X band receivers) from the GBT archive. We implement a low-complexity statistical process to vet out RFI and highlight signals that, upon visual inspection, appear more promising than those from previous analyses. Our work returns candidate signals found previously using both traditional and machine learning algorithms, as well as many promising ones not previously identified. This analysis represents the largest dataset searched for technosignatures to date, and highlights the efficacy that traditional (non-machine-learning) algorithms continue to have in these types of technosignature searches. We find that less than 1% of stars host transmitters brighter than $\sim 0.3$ Arecibo radar equivalents broadcasting in our direction over the frequency band covered.
@@ 1 Introduction
The largest program currently driving the field of SETI research is the Breakthrough Listen (BL) Initiative (Worden et al., 2017), which has as part of its core mission a survey of 1 million nearby stars. Over the past nine years, BL has been using the 100-m Robert C. Byrd Green Bank Telescope (GBT) in West Virginia to survey thousands of stars (MacMahon et al. 2018, Price et al. 2020), close to a hundred galaxies (Choza et al., 2024), and a variety of solar system objects (Enriquez et al., 2018) at frequencies of 1–12 GHz. BL also observes at other facilities around the globe, including the Murriyang telescope at Parkes Telescope in Australia (Price et al., 2018). BL’s projects on GBT and Parkes have resulted in the aggregation of over 20 PB of radio data.
@@
While turboSETI has helped provide some of the best constraints on extraterrestrial life to date, it has some limitations. Some signals (including some narrowband Doppler drifting signals, as well as the broader class of signals with other morphologies) are missed by turboSETI (Ma et al., 2023). Additionally, in each GBT cadence turboSETI typically finds thousands of hits, and several events (Choza et al., 2024), requiring tedious visual inspection. In response, other authors have come up with their own algorithms to find higher-quality signals, including machine learning approaches (Ma et al., 2023; Brzycki et al., 2020; Pinchuk & Margot, 2022; Gajjar et al., 2022; Jacobson-Bell et al., 2025). Some examples are the use of convolutional neural networks and random forest decision trees to recognize and classify specific kinds of data, as well as generative adversarial networks to pick out anomalies. These have proven successful in handling the large datasets involved in technosignature searches and have picked up promising candidates, but come with their own respective problems. Specifically, these are the high computational costs required to train the neural networks, the dependence on training data, and the lack of interpretability in the decision-making process. Deep learning models also only make predictions based on the data they have been trained on, and cannot extrapolate to novel signal types that may not have been encompassed in the training set. This is one advantage of an unsupervised learning approach (see Jacobsen-Bell 2024 for example).
@@ 2 Dataset
We deployed our search algorithm on the entire GBT dataset in the Breakthrough Listen data center at the University of California, Berkeley as of 2023 July; we did not analyze additional GBT data stored onsite at the Green Bank Observatory. This comprises 4,477 hours of of on-sky data across 9,684 cadences and 3350 unique pointings (of which 3077 are stars, and the rest are galaxies, globular clusters, and solar system objects). The positions of each source is shown in Figure 1, along with a histogram of all their distances in Figure 2. This makes this study one of the largest SETI search to date in terms of number of targets and observations. The observations come from hundreds of different observing sessions over the past eight years. The dataset used covers all four of the main bands used to date in BL’s GBT searches, namely the 1.1–1.9 GHz L band, the 1.8–2.7 GHz S band, the 4.0–7.8 GHz C band, and the 7.8–11.2 GHz X band. The full number of cadences and unique targets in each band can be found in Table 1. Each cadence in this dataset is composed of six 5-minute observations, three on-source and three off-source. The data is recorded in HDF5 format, with a time resolution of $\sim 18$ s and a fine frequency resolution of 2.79 Hz for 99.3% of the data and 2.84 Hz for the other 0.7% (Lebofsky et al., 2019).
Each BL compute node captures a portion of the total bandwidth during observing with a maximum of 187.5 MHz. Of the 9,684 cadences in our sample, 93.9% have been stored as individual 187.5 MHz data products. These have been grouped with the other five 187.5 MHz data products spanning the same frequency range, but corresponding to the other ON and OFF scans in the cadence. These groups of six 187.5 MHz data products form what we label a ‘sub-cadence’, and we split these sub-cadences into subsamples that we process in parallel. 6.1% of our observations have been spliced together in frequency into one large data product spanning the full band (as much as 4 GHz). Due to the differences in size, we combine all of these spliced data products into their own sub-sample. This leaves us with 80,740 sub-cadences, which we group into 81 sub-samples each containing 1,000 sub-cadences (except for the last subsample which contains slightly fewer).
@@
While all cadences used alternate between ON-source and OFF-source observations, two different patterns are used. Earlier cadences (prior to 2016 April 26) follow an ABABAB source pattern, with the same offset position (B) being used on all three OFF scans between the primary source (A) scans. All other cadences use an ABACAD pattern, with a different source (B,C,D) used in each OFF scan. For the ABABAB observations, the angular offset between ON and OFF is 2° in declination. For the ABACAD observations, the secondary targets are at least five beam-widths away from the primary target. For the purposes of our search we treat both type of cadence scheme equally.
Overall, 534 TB of data has been analyzed in this work. Some of the data analyzed was included in previously published work (see Section 5.5), so we are also able to benchmark the results from this search against earlier ones. This gives some reference for our algorithm’s performance against both other classical algorithms (such as the standard de-Doppler technique used by turboSETI), as well as machine learning techniques.
@@ 3.2 Algorithm
The novel algorithm we present in this study is dubbed pickles [^2] https://github.com/UCBerkeleySETI/pickles, Pipeline for Identification of Candidates using Kurtosis to Locate Extraterrestrial Signals. The core idea is that a narrow subset of the frequency space without any signal present will be almost entirely Gaussian noise, while such a frequency subset with a signal present will have a power spectrum that deviates significantly from a Gaussian distribution. We label these subsets of the frequency space frequency ‘blocks’, and they correspond to a sliver of a cadence’s frequency coverage. They consist of a set number of frequency channels from all six observations in a cadence, all centered on the same frequency.
@@
where $\sigma$ is the standard deviation of the spectra in the frequency block, $x_{i}$ are the individual intensity values, $\bar{x}$ the mean intensity, and N the number of data points. On the other hand, a frequency block with a signal in it could have a much higher kurtosis (we will define what ‘high’ could mean in Section 3.3). We note that this use of kurtosis is different than previous uses of kurtosis for RFI flagging Nita et al. (2007), where a ‘spectral’ kurtosis is instead calculated on the time series of each channel, yielding multiple kurtosis values. Here, we instead compute a ‘block’ kurtosis, where we flatten our entire frequency block across multiple time channels into a 1D dataset. We refer to this ‘block’ kurtosis as kurtosis for simplicity.
@@
A promising ET candidate signal would manifest only in the ON scans and not in the OFF scans of a cadence. The frequency block this signal falls in would thus have a high kurtosis in all of its ON scans and a very low kurtosis in all of its OFF scans. By dividing up the cadence’s full frequency range into narrow frequency blocks, we can thus test each block against this condition. While the requirement for a low kurtosis in the OFF observations means that any genuine ET signal in the same frequency block as RFI will not be found, this is a trade-off we accept in exchange for keeping our algorithm low in complexity and reducing the number of false positives.
@@
Freq Double Frequency in the observation where a candidate was found.
@@
The size of the frequency block that we split each cadence into is an important choice. A smaller block will produce a higher kurtosis compared to the same signal in a larger block, which enables the detection of weaker signals. It also reduces the probability that RFI falls into the same block as a genuine signal. However, if the block is too small then a high-drift signal might only appear in one of the ON scans. We opt for a frequency block of 1024 fine channels, equivalent to $2.86$ kHz, dividing up each cadence into $n=N_{\rm chan}/1024$ separate blocks where $N_{\rm chan}$ is the total number of fine channels in the file. We then create an additional $n-1$ frequency blocks of the same size, each centered on the divide between the original regions. This gives a total of $12n-$ 1 overlapping frequency blocks. We then remove all frequency blocks falling into the high RFI regions and notch-filter regions defined above.
Loading arrays of length power of 2 into memory uses this available memory better as the hardware and the operating system manage memory in units of power of two. As a result we initially consider four potential block sizes, of 512, 1024, 2048, and 4096 channels. We opt to divide our observations into frequency blocks of 1024 fine channels for two main reasons:
  1. 1.
    Figure 4: Kurtosis as a function of signal strength for a narrowband signal present in each time bin of an observation. Only frequency blocks of size 512 and 1024 have an expected kurtosis $>1$ for a signal with SNR=10.
  2. 2.
    We want the size of our blocks to capture drifting signals. A signal might drift out of frame if a block size is too small. Figure 5 shows the probability of a signal being present in at least 2 of the 3 ON scans of a cadence for blocks of varying size, as a function of the drift rate of a signal. We aim to be sensitive to signals drifting up to ±$\pm 4$ Hz/s, consistent with past Breakthrough Listen searches. The 512 channel block has a 0% probability of capturing a signal drifting $\geq 3$ Hz/s. Blocks of size 1024 channels have a 98.7% chance of capturing some portion of a signal drifting at ±$\pm 4$ Hz/s in at least two of the three ON scans. We thus choose to use blocks of size 1024 channels, as they have a near 100% chance of capturing signals drifting ±$\pm 4$ Hz/s and will produce kurtoses $>1$ if they contain signals with SNR $\approx$ 10. We define the SNR of a signal as the signal amplitude divided by the median absolute deviation of the noise in its block.
The next step is to remove all blocks that we believe are unlikely to contain a signal. Since input/output I/O read time is the most time-intensive part of this algorithm, we would like to avoid analyzing blocks that contain only noise in all of their ON observations. To do this we load the entire frequency data of one time slice in each ON observation (the choice of time slice is somewhat arbitrary, we use the eighth), and then split this array into the frequency blocks that were created earlier. This results in a set of length 1024 arrays, each representing one time slice of a 2.86 kHz width block, which are each individually checked to see if they contain any peaks $>3\sigma$. Those that do are checked again to see if they contain any peaks with a median absolute deviation $>5\sigma$, and the frequency blocks that do are recorded as ‘hotspots’. This is repeated for the equivalent time slices of the second and third ON scans, with all hotspots recorded for each. All frequency blocks that contain at least two ON observations labeled as hotspots are deemed likely to have a signal in at least two of the ON observations, and are kept for full analysis. These are deemed Level 0 frequency blocks, and get passed on to the rest of our pipeline for full analysis.
The most time intensive step in our algorithm is the I/O time of loading individual blocks of an observation from the BL data archive into memory. To speed this up, we divide each full frequency band into 8 sub-bands. We load the sub-bands of each observation in a cadence into memory at once, and analyze all frequency blocks present in that sub-band before moving on to the next sub-band. This allows us to load observations faster into memory compared to reading in and processing each 1024-channel snippet one by one, because of the switching cost during (I/O) vs processing.
Table 4: High-Density RFI Removed Frequencies
Frequency Federal Allocation
$(\mathrm{MHz})$
$1164-1200$ Global Navigation Satellite System (GNSS)
$1200-1340$ Notch Filter
$1370-1390$ Air traffic control (ATC)
$1520-1560$ Mobile-satellite service (MSS)
$1560-1630$ GNSS
$1670-1705$ Geostationary operational
environmental satellite (GOES)
$1915-2000$ Personal communications services (PCS)
$2025-2035$ GOES
$2100-2120$ NASA Deep Space Network use (DSN)
$2180-2200$ MSS
$2200-2280$ Earth exploration satellite (EES)
$2300-2360$ Notch Filter
$2485-2500$ MSS
$4680-4800$ Fixed Satellite Service (FSS)
$8150-8350$ EES
$9550-9650$ EES
$10700-11200$ FSS
@@ 3.3 Filtering
The properties recorded above are used to select good candidates from the set of frequency blocks analyzed. The kurtosis statistic is the primary metric. The goal is to find frequency blocks with high kurtoses in the ON scans, and low kurtoses in the OFF scans. We impose a cutoff of $k_{\rm ON,med}\geq 0.5$, where $k_{\rm ON,med}$ is the median of the three ON scan kurtoses. This is to take into account that a rapidly drifting signal (close to $\pm$4 Hz/s) may not appear in all 3 ON observations, but should still appear in at least 2. The median kurtosis value of the ON scans should thus be $>1$ if a strong (SNR $>10$) signal is present. For the block size we use here, a narrowband signal with a SNR of 10 would produce an associated kurtosis of $\gtrsim 1$ when injected into a Gaussian noise background (see Section 3.2). In theory this means we are sensitive to signals with even lower SNR, closer to SNR $\approx$ 8. We require that the kurtosis of each OFF scan is $\leq 0.25$, so that there is little or no signal in each OFF scan. This $k_{\rm OFF,max}$ cutoff can be lowered to decrease the number of false positives.
After imposing the kurtosis cuts detailed above, we are left with $170,623$ ‘candidate’ frequency blocks from the original $23,688,878$ analyzed. This is about a $0.7\%$ detection rate, or equivalently a $0.7\%$ false positive rate if we assume essentially all detected signals are RFI. These are still too many candidates to inspect visually, so additional filters must be applied. We use two: a requirement that any signal present in the frequency block have non-zero drift; and that the signal present is not a broadband streak present in only a small number of time bins.
Our drift-filter checks a signal’s drift rate by comparing the location of the signal in each ON scan. We integrate each ON scan along the time dimension. A signal with zero drift will appear in the same frequency bin in each time slice, and thus show up clearly when the ON scan is time-integrated. We record the frequency location of these signals by running a peak-finding algorithm on the time-integrated ON scans, and compare the frequency locations of the peaks between each ON scan using the Python package scipy’s findpeaks routine. If the peaks in adjacent ON scans are less than 2 frequency bins apart (indicating a drift rate of $<0.01$ Hz/s) then we mark the signal as having zero drift.
@@
Our broadband filter functions by checking if there are large changes in intensity over time in an observation. Broadband interference often presents itself as horizontal streaks across the entire frequency range of a block, but confined to just one or two time bins. To identify such contamination, we average across the frequency of a block’s ON scan. The broadband interference will show up as a spike in intensity at a certain time over the course of the observation. If the peak has a median absolute deviation $\geq 10$ from baseline, we classify it as likely being produced by broadband streaks and reject the signal. Of course, we might also discard some real pulsed technosignatures using this method, but only if they have pernicious duty cycles that happen to line up with the duration of our ON observations. In general, these pulsed signals belong to a different kind of transmitter than this search is optimized for, as our data has a limited time resolution of $\sim 18$s. There have been other search approaches built to target pulse beacons (Gajjar et al. 2022, Suresh et al. 2023). Our broadband filter is illustrated in Figure 6.
A final common false positive is small ‘blips’ of high intensity, which show up in a small number of frequency and time bins in an otherwise empty block. If such blips show up in two or more ON scans and are high enough intensity to significantly skew the kurtosis, then their respective blocks will pass our kurtosis filter. Since these small blips only occupy a few frequency bins, they are averaged out when the entire scan is frequency integrated, and our broadband filter does not pick them up. To reduce the number of such false positives that make it through our pipeline we calculate the maximum of each time slice, and check if there are significant outliers, defined as $>10\sigma$, within these local maxima. Signals that are fairly constant in intensity over time will not be eliminated. However, this method may eliminate true signals that have a significant intensity spike in one time bin.

### v2

@@ Abstract
The Breakthrough Listen program is, to date, the most extensive search for technological life beyond Earth. Over the past nine years it has surveyed thousands of nearby stars and close to 100 nearby galaxies with telescopes around the world, including the Robert C. Byrd Green Bank Telescope (GBT) in West Virginia. The goal is to find evidence of technosignatures of other civilizations, such as narrowband Doppler drifting radio signals. Despite the GBT’s location in a radio-quiet zone, the primary challenge of this search continues to be the ability to pick out genuine candidates from the high quantities of human-generated radio-frequency interference (RFI). Here we present a novel search method aimed at finding these ‘needle-in-a-haystack’ type signals, applied to 6,630 observation cadences of 2,623 stars (each observed with one or more of the L, S, C, and X band receivers) from the GBT archive. We implement a low-complexity statistical process to vet out RFI and highlight signals that, upon visual inspection, are less evidently RFI than those from previous analyses. Our work returns candidate signals found previously using both traditional and machine learning algorithms, as well as many not previously identified. This analysis represents the largest dataset searched for technosignatures to date, and highlights the efficacy that traditional algorithms continue to have in these types of technosignature searches. We find that less than 1% of stars host transmitters brighter than $\sim 0.3$ Arecibo radar equivalents broadcasting in our direction over the frequency band covered.
Radio Astronomy — Search for Extraterrestrial Intelligence — Technosignatures
@@ 1 Introduction
The largest program currently driving the field of SETI research is the Breakthrough Listen (BL) Initiative (Worden et al., 2017), which has as part of its core mission a survey of 1 million nearby stars. Over the past nine years, BL has been using the 100-m Robert C. Byrd Green Bank Telescope (GBT) in West Virginia to survey thousands of stars (MacMahon et al. 2018, Price et al. 2020), close to a hundred galaxies (Choza et al., 2024), and a variety of solar system objects (Enriquez et al., 2018) at frequencies of 1–12 GHz. BL also observes at other facilities around the globe, including the Murriyang telescope at Parkes in Australia (Price et al., 2018). BL’s projects on GBT and Parkes have resulted in the aggregation of over 20 PB of radio data.
@@
While turboSETI has helped provide some of the best constraints on extraterrestrial life to date, it has some limitations. Some signals (including some narrowband Doppler drifting signals, as well as the broader class of signals with other morphologies) are missed by turboSETI (Ma et al., 2023). Additionally, in each GBT cadence turboSETI typically finds thousands of hits (signals detected by turboSETI’s narrowband Doppler drifting search), and several events (Choza et al., 2024), requiring tedious visual inspection. In response, other authors have come up with their own algorithms to find higher-quality signals, including both traditional (Margot et al., 2023) as well as machine learning approaches (Ma et al., 2023; Brzycki et al., 2020; Pinchuk & Margot, 2022; Gajjar et al., 2022; Jacobson-Bell et al., 2025a). Some examples of machine learning approaches include the use of convolutional neural networks and random forest decision trees to recognize and classify specific kinds of data, as well as generative adversarial networks to pick out anomalies. These have proven successful in handling the large datasets involved in technosignature searches, but come with their own respective problems. Specifically, these are the high computational costs required to train the neural networks, the dependence on training data, and the lack of interpretability in the decision-making process. Deep learning models also only make predictions based on the data they have been trained on, and cannot extrapolate to novel signal types that may not have been encompassed in the training set. This is one advantage of an unsupervised learning approach (see Jacobson-Bell et al. 2025b for example).
@@ 2 Dataset
We deployed our search algorithm on a large fraction of the entire GBT dataset in the Breakthrough Listen data center at the University of California, Berkeley as of 2023 July; we did not analyze additional GBT data stored onsite at the Green Bank Observatory. This comprises 3,165 hours of of on-sky data across 6,630 cadences and 2,841 unique pointings (of which 2,623 are stars, and the rest are galaxies, globular clusters, and solar system objects) [^2] The sample size has been corrected from the original version of this paper, and the affected figures, tables, and transmitter rates have been updated.. The position of each source is shown in Figure 1, along with a histogram of all their distances in Figure 2. This makes this study one of the largest SETI searches to date in terms of number of targets and observations. The observations come from hundreds of different observing sessions over the past eight years. The dataset used covers all four of the main bands used to date in BL’s GBT searches, namely the 1.1–1.9 GHz L band, the 1.8–2.7 GHz S band, the 4.0–7.8 GHz C band, and the 7.8–11.2 GHz X band. The full number of cadences and unique targets in each band can be found in Table 1. Each cadence in this dataset is composed of six 5-minute observations, three on-source and three off-source. These data are recorded in HDF5 format, with a time resolution of $\sim 18$ s and a fine frequency resolution of 2.79 Hz for 96.4% of the data and 2.84 Hz for the other 3.6% (Lebofsky et al., 2019).
Each BL compute node captures a portion of the total bandwidth during observing with a maximum of 187.5 MHz. Of the 6,630 cadences in our sample, 67.8% have been stored as individual 187.5 MHz data products. These have been grouped with the other five 187.5 MHz data products spanning the same frequency range, but corresponding to the other ON and OFF scans in the cadence. These groups of six 187.5 MHz data products form what we label a ‘sub-cadence’, and we split these sub-cadences into subsamples that we process in parallel. The other 32.2% of our observations have been spliced together in frequency into one large data product spanning the full band (as much as 4 GHz). Due to the differences in size, we combine all of these spliced data products into their own sub-sample. This leaves us with 62,308 sub-cadences, which we group into 63 sub-samples each containing 1,000 sub-cadences (except for the last subsample which contains slightly fewer).
@@
While all cadences used alternate between ON-source and OFF-source observations, two different patterns are used. Earlier cadences (prior to 2016 April 26) follow an ABABAB source pattern, with the same offset position (B) being used on all three OFF scans between the primary source (A) scans. All other cadences use an ABACAD pattern, with a different source (B,C,D) used in each OFF scan. For the ABABAB observations, the angular offset between ON and OFF is 2° in declination (for comparison, the GBT FWHM is $11\farcm 5$ at 1.1 GHz). For the ABACAD observations, the secondary targets are at least five beam-widths away from the primary target. For the purposes of our search we treat both type of cadence scheme equally.
Some of the data analyzed were included in previously published work (see Section 5.5), so we are also able to benchmark the results from this search against earlier ones. This gives some reference for our algorithm’s performance against both other classical algorithms (such as the standard de-Doppler technique used by turboSETI), as well as machine learning techniques.
@@ 3.2 Algorithm
The novel algorithm we present in this study is dubbed pickles [^3] https://github.com/UCBerkeleySETI/pickles, Pipeline for Identification of Candidates using Kurtosis to Locate Extraterrestrial Signals. The core idea is that a narrow subset of the frequency space without any signal present will be almost entirely noise, while such a frequency subset with a signal present will have a power spectrum that deviates significantly from the background noise distribution. We label these subsets of the frequency space as ‘frequency blocks’ or ‘blocks’ for short, and they correspond to a sliver of a cadence’s frequency coverage. They consist of a set number of frequency channels from all six observations in a cadence, all centered on the same frequency.
@@
where $\sigma$ is the standard deviation of the spectra in the frequency block, $x_{i}$ are the individual intensity values, $\bar{x}$ the mean intensity, and N the number of data points. On the other hand, a frequency block with a signal in it could have a much higher kurtosis (we will define what ‘high’ could mean in Section 3.3). We note that this use of kurtosis is different than previous uses of kurtosis for RFI flagging (Nita et al., 2007), where a ‘spectral’ kurtosis is instead calculated on the time series of each channel, yielding multiple kurtosis values. Here, we instead compute a ‘block’ kurtosis, where we flatten our entire frequency block across multiple time channels into a 1D dataset. We refer to this ‘block’ kurtosis as kurtosis for simplicity.
@@
A strong candidate ET signal would manifest only in the ON scans and not in the OFF scans of a cadence. The frequency block this signal falls in would thus have a high kurtosis in all of its ON scans and a very low kurtosis in all of its OFF scans. By dividing up the cadence’s full frequency range into narrow frequency blocks, we can thus test each block against this condition. While the requirement for a low kurtosis in the OFF observations means that any genuine ET signal in the same frequency block as RFI will not be found, this is a trade-off we accept in exchange for keeping our algorithm low in complexity and reducing the number of false positives.
The size of the frequency block that we split each cadence into is an important choice. A smaller block will produce a higher kurtosis compared to the same signal in a larger block, which enables the detection of weaker signals. It also reduces the probability that RFI falls into the same block as a genuine signal. However, if the block is too small then a high-drift signal might only appear in one of the ON scans. We opt for a frequency block of 1024 fine channels, equivalent to $2.86$ kHz, dividing up each cadence into $n=N_{\rm chan}/1024$ separate blocks where $N_{\rm chan}$ is the total number of fine channels in the file. We then create an additional n1n-1 frequency blocks of the same size, each centered on the divide between the original regions. This gives a total of 2n12n-1 overlapping frequency blocks. We then remove all frequency blocks falling into the high RFI regions and notch-filter regions defined above.
Loading arrays of length power of 2 into memory uses this available memory better as the hardware and the operating system manage memory in units of power of two. As a result we initially considered four potential block sizes, of 512, 1024, 2048, and 4096 channels. We opt to divide our observations into frequency blocks of 1024 fine channels for two main reasons:
  1. 1.
    Figure 4: Kurtosis as a function of signal strength for a narrowband signal present in each time bin of an observation. Only frequency blocks of size 512 and 1024 have an expected kurtosis >1>1 for a signal with SNR=10. This is true for L, S, C, and X bands.
  2. 2.
    We want the size of our blocks to capture drifting signals. A signal might drift out of frame if a block size is too small. Figure 5 shows the probability of a signal being present in at least 2 of the 3 ON scans of a cadence for blocks of varying size, as a function of the drift rate of a signal. We aim to be sensitive to signals drifting up to $\pm 4$ Hz/s, consistent with past Breakthrough Listen searches. The 512 channel block has a 0% probability of capturing a signal drifting $\geq 3$ Hz/s. Blocks of size 1024 channels have a 98.7% chance of capturing some portion of a signal drifting at $\pm 4$ Hz/s in at least two of the three ON scans. We thus choose to use blocks of size 1024 channels, as they have a near 100% chance of capturing signals drifting $\pm 4$ Hz/s and will produce kurtoses >1>1 if they contain signals with SNR $\approx$ 10. We define the SNR of a signal as the signal amplitude divided by the median absolute deviation of the noise in its block.
The next step is to remove all blocks that we believe are unlikely to contain a signal. Since input/output (I/O) read time is the most time-intensive part of this algorithm, we would like to avoid analyzing blocks that contain only noise in all of their ON observations. To do this we load the entire frequency data of one time slice in each ON observation (the choice of time slice is somewhat arbitrary, we use the eighth), and then split this array into the frequency blocks that were created earlier. This results in a set of length 1024 arrays, each representing one time slice of a 2.86 kHz frequency block, which are each individually checked to see if they contain any peaks with a median absolute deviation $5\sigma$, and the frequency blocks that do are recorded as ‘hotspots’. This is repeated for the equivalent time slices of the second and third ON scans, with all hotspots recorded for each. All frequency blocks that contain at least two ON observations labeled as hotspots are deemed likely to have a signal in at least two of the ON observations, and are kept for full analysis. These are deemed Level 0 frequency blocks, and get passed on to the rest of our pipeline for full analysis (see Section 3.4 for more details).
The most time intensive step in our algorithm is the I/O time of loading individual blocks of an observation from the BL data archive into memory. To speed this up, we divide each full frequency band into 8 sub-bands. We load the sub-bands of each observation in a cadence into memory at once, and analyze all frequency blocks present in that sub-band before moving on to the next sub-band. This allows us to load observations faster into memory compared to reading in and processing each 1024-channel frequency block one by one, because of the switching cost during I/O vs processing.
Table 3: High-Density RFI Removed Frequencies
Frequency Federal Allocation
$(\mathrm{MHz})$
116412001164-1200 Global Navigation Satellite System (GNSS)
120013401200-1340 Notch Filter
137013901370-1390 Air traffic control (ATC)
152015601520-1560 Mobile-satellite service (MSS)
156016301560-1630 GNSS
167017051670-1705 Geostationary operational
environmental satellite (GOES)
191520001915-2000 Personal communications services (PCS)
202520352025-2035 GOES
210021202100-2120 NASA Deep Space Network use (DSN)
218022002180-2200 MSS
220022802200-2280 Earth exploration satellite (EES)
230023602300-2360 Notch Filter
248525002485-2500 MSS
468048004680-4800 Fixed Satellite Service (FSS)
815083508150-8350 EES
955096509550-9650 EES
107001120010700-11200 FSS
@@
Freq Double Frequency in the observation where the signal was found.
@@ 3.3 Filtering
The properties recorded above are used to select strong candidates from the set of frequency blocks analyzed. The kurtosis statistic is the primary metric. The goal is to find frequency blocks with high kurtoses in the ON scans, and low kurtoses in the OFF scans. We impose a cutoff of $k_{\rm ON,med}\geq 0.5$, where $k_{\rm ON,med}$ is the median of the three ON scan kurtoses. This is to take into account that a rapidly drifting signal (close to $\pm$4 Hz/s) may not appear in all 3 ON observations, but should still appear in at least 2. The median kurtosis value of the ON scans should thus be >1>1 if a strong (SNR >10>10) signal is present. For the block size we use here, a narrowband signal with a SNR of 10 would produce an associated kurtosis of $\gtrsim 1$ when injected into a background of sampled GBT receiver noise (see Section 3.2). In theory this means we are sensitive to signals with even lower SNR, closer to SNR $\approx$ 8. We require that the kurtosis of each OFF scan is $\leq 0.25$, so that there is little or no signal in each OFF scan. This $k_{\rm OFF,max}$ cutoff can be lowered to decrease the number of false positives.
After imposing the kurtosis cuts detailed above, we are left with $175,160$ frequency blocks from the original $24,368,715$ analyzed. This is about a $0.7\%$ detection rate, or equivalently a $0.7\%$ false positive rate if we assume essentially all detected signals are RFI. These are still too many frequency blocks to inspect visually, so additional filters must be applied. We use two: a requirement that any signal present in the frequency block have non-zero drift; and that the signal present is not a broadband streak present in only a small number of time bins.
Our drift-filter checks a signal’s drift rate by comparing the location of the signal in each ON scan. We integrate each ON scan along the time dimension. A signal with zero drift will appear in the same frequency bin in each time slice, and thus show up clearly when the ON scan is time-integrated. We record the frequency location of these signals by running a peak-finding algorithm on the time-integrated ON scans, and compare the frequency locations of the peaks between each ON scan using the Python package scipy’s findpeaks routine. If the peaks in adjacent ON scans are less than 2 frequency bins apart (indicating a drift rate of <0.01<0.01 Hz/s) then we mark the signal as having zero drift.
@@
Our broadband filter functions by checking if there are large changes in intensity over time in an observation. Broadband interference often presents itself as horizontal streaks across the entire frequency range of a block, but confined to just one or two time bins. To identify such contamination, we average across the frequency of a block’s ON scan. The broadband interference will show up as a spike in intensity at a certain time over the course of the observation. If the peak has a median absolute deviation $\geq 10$ from baseline, we classify it as likely being produced by broadband streaks and reject the signal. Of course, we might also discard some real pulsed technosignatures using this method, but only if they have pernicious duty cycles that happen to line up with the duration of our ON observations. In general, these pulsed signals belong to a different kind of transmitter than this search is optimized for, as our data have a limited time resolution of $\sim 18$s. There have been other search approaches built to target pulse beacons (Gajjar et al. 2022, Suresh et al. 2023). Our broadband filter is illustrated in Figure 6.
A final common false positive is small ‘blips’ of high intensity, which show up in a small number of frequency and time bins in an otherwise empty block. If such blips show up in two or more ON scans and are high enough intensity to significantly skew the kurtosis, then their respective blocks will pass our kurtosis filter. Since these small blips only occupy a few frequency bins, they are averaged out when the entire scan is frequency integrated, and our broadband filter does not pick them up. To reduce the number of such false positives that make it through our pipeline we calculate the maximum of each time slice, and check if there are significant outliers, defined as $10\sigma$, within these local maxima. Signals that are fairly constant in intensity over time will not be eliminated. However, this method may eliminate true signals that have a significant intensity spike in one time bin.